IODP Proceedings    Volume contents     Search



Shimizu, Japan, port call

Expedition 338 officially began at 0000 h on 1 October 2012 while the Chikyu was north of Izu-Oshima, Japan, as part of the evacuation procedures for Typhoon Jelawat. The evacuation ended at 0400 h on 1 October when the vessel began to return to port call in Shimizu, Japan. Port call began at 0550 h on 2 October, with all loading operations concluded by 2400 h on 3 October. The vessel departed Shimizu, Japan, for Site C0002 at 1035 h on 4 October.

Site C0002

Hole C0002F

Seabed survey and transponder deployment began at 1230 h on 5 October 2012 and were completed on 7 October. After calibrating transponders, the Chikyu sailed to the blowout preventer (BOP) and riser running point 20.8 nmi northwest of Site C0002. The BOP and riser joints were set up in the moonpool and prepared for running, which began at 0400 h on 8 October. By 0400 h on 9 October, the BOP reached 490 m drilling depth below rig floor (DRF); however, because of the development of Typhoon Prapiroon, a decision was made to wait on weather (WOW) and to monitor the typhoon track at 0900 h on 9 October; subsequently, BOP and riser recovery began for evacuation standby at 1300 h on 11 October. Recovery was completed at 0600 h on 12 October. The vessel moved 21 nmi northwest of Site C0002 to facilitate the arrival of the first scientists by helicopter transfer on 13 October and to avoid the strengthening current. We continued to monitor developing typhoon tracks and remained in WOW status until 0700 h on 19 October, when preparations to begin running the BOP into the moonpool commenced. Once the BOP was in the moonpool at ~1700 h, BOP pressure tests were run, ending at 1800 h, and the BOP was run into the water. Careful attention was paid to the auxiliary (AUX) line and buoyant riser joints, all the while conducting regular AUX pressure tests (six pressure tests in total). AUX line troubleshooting and buoyancy riser joint replacement was required on 21 October when one joint was replaced.

While running the BOP, several function tests and pressure tests found failures on the AUX line, the conduit line, and the hot line, which were subsequently repaired. The BOP finally landed on the wellhead at 2130 h on 26 October. All function tests, repairs, and maintenance were complete by 1045 h on 29 October. The tests included a pickup test by increasing tensioner tension, slump tests by decreasing drawworks tension, function tests of the diverter, pressure tests of the wellhead connector and the 20 inch casing, function tests of the BOP from both the Blue Pod (driller’s control panel) and the Yellow Pod (toolpusher’s control panel), a pressure test of the BOP, a pressure test of the BOP with 20 inch casing and inside drill pipe, a function test of the remaining valves in block position, and confirmation of working time and flow rate of the BOP.

The bottom-hole assembly (BHA) for drill-out cement (DOC) was run into the hole at 1900 h on 30 October and tagged the top of the cement/bottom of the hole (842 mbsf) at 1415 h on 31 October. DOC with the 17 inch bit reached 872.5 mbsf, with an extension into the formation (to 875.5 mbsf) to confirm the cement plug was completely drilled through that ended at 2045 h. After DOC was complete, the hole was swept with Hi-Vis mud, then seawater, and then KNPP mud. Methane (15.5%) was found in the drilling mud, and after monitoring, it was decided to move ahead with two leak-off tests (LOTs) of the 20 inch casing shoe starting at ~1100 h on 1 November.

The DOC BHA was recovered and laid down, after which the LWD/measurement-while-drilling (MWD) underreamer BHA (see “Introduction” in the “Methods” chapter [Strasser et al., 2014a]) was made up and run into the hole (Table T1) and tested at 0430 h on 2 November. All function tests (n = 4: shallow, 50 meters below sea level [mbsl], and 132.5 and 778.5 mbsf) were successfully completed by 0100 h on 3 November, after which drilling began. The BHA was picked up to 905.8 mbsf after drilling ahead to 915.19 mbsf at 0745 h on 4 November to activate the underreamer (total of six attempts made), which took ~2 h.

Drilling with the underreamer BHA began at 1000 h on 4 November. Rate of penetration (ROP) was controlled at 40 m/h from 914.5 mbsf and changed to 20–26 m/h below 923.8 mbsf. Concerns with the large volume of cuttings, more than the waste mud control system could handle, required careful control, sweeping, and sometimes suspension of drilling. Repeat logging was carried out three times using these periods (~30 m/h): uplogging from 1432.40 to 1494.25 mbsf, downlogging from 1480.86 to 1538.77 mbsf, and downlogging from 1557.82 to 1615.58 mbsf. Expected rough weather caused drilling to be suspended at 1538.5 mbsf (2200 h on 5 November), and the BHA was pulled out of the hole above the BOP. Drilling resumed (ROP = 15 m/h) at 2100 h on 6 November but stopped again when 1604.5 mbsf was reached because dispersed sediment was observed coming from one port on the 36 inch conductor pipe with subsequent hydrate accumulation around the wellhead (1230 h on 7 November). Observation for 6 h showed stable flow, so drilling was resumed (ROP = 10–15 m/h) at 2300 h on 7 November. From 1604.5 to 2005.5 mbsf (2300 h on 7 November to 0800 h on 11 November), drilling continued with some interruptions, mainly because of waste mud control system issues (e.g., mud pump problems, screw conveyer for mud transfer issues, and strong currents interfering with supply boats). Drilling was stopped again to standby for rough weather, and the BHA was pulled out of the hole above the BOP at 1515 h on 12 November.

During WOW, a critical function failure of the Double V Shear (DVS) ram (a key component of the BOP designed to close the hole while maintaining annular pressure after the BOP disconnects) was found. It was decided at 0530 h on 13 November to spot cement and suspend the hole until the failure was resolved. Fortunately, the root cause of the failure was determined, and a mitigation plan was enacted for the BOP; therefore, at 0230 h on 14 November, the BOP disconnection plan was canceled. DOC began at 0115 h on 15 November and completed at 1630 h when the DOC BHA was pulled out of the hole to the surface. Subsequently, the underreamer BHA for reaming and enlarging the hole from 12¼ to 20 inches was run into the hole at 0000 h on 16 November. The underreamer was activated at 2848 m DRF at 0300 h; reaming was conducted from 0300 to 2300 h on 16 November. When bit depth reached 3955 m DRF (1990 mbsf) and the underreamer depth and 20 inch hole was at 3949 m DRF (1984 mbsf), an approaching weather system caused another WOW. Reaming and backreaming down to the final bit depth were performed over the next 5 h until the BHA was pulled out of the hole to 1576 m DRF by 1430 h on 17 November after performing backreaming in the hole and spotting with high-viscosity mud. As a standby operation, the kill line was flushed and the DVS ram was closed.

At 1730 h on 17 November, the vessel went into an emergency disconnect because of the high current speed of ~4.5 kt from the west and a sudden change in wind direction and speed (Fig. F5), the combination of which forced the vessel 40 m from the well center, and control of the vessel was lost. The emergency disconnect was conducted safely and efficiently and then the vessel drifted 1300 m east of the wellhead at 4 kt. However, while this was happening, the riser pipe hit the hull of the vessel and the intermediate flex joint sustained damage because of significant flexure. Pulling out of the hole to the surface continued, ending at 0415 h on 18 November; whereupon the vessel recovered the remotely operated vehicle (ROV) and moved to the low-current area (LCA) 13 nmi northwest of Hole C0002F to begin riser recovery. The tensioner ring was removed, and electrical cable connections were also removed.

Riser pull out began at 2300 h and termination joints were checked. The diverter was laid out from 0000 h on 19 November, and the riser termination joint was pulled out of the moonpool at 0430 h. The hot line was reterminated in the moonpool, and a pressure check from 1400 to 1800 h confirmed the BOP status was still good. The gooseneck was inspected, including welding points on the slip joint with magnetic particle inspection (MPI) for cracks from 2100 to 0030 h on 20 November. Riser sections continued to be pulled, stopping to perform pressure tests on the choke and kill lines for 3 h from 0315 h. After the intermediate flex joint was pulled up, the termination joint weld was inspected via MPI, whereupon four cracks were found; advice on repair was sought from the manufacturer. The results of careful inspection and manufacturer advice led to the conclusion that the repairs needed could only be conducted on shore; accordingly, it was decided to end riser operations as the time needed for repairs would exceed the planned expedition duration.

Meanwhile, the conduit and kill lines on the gooseneck passed inspection by 2315 h on 21 November after several tests and leakage troubleshooting. Once passed, the gooseneck was removed so that the intermediate flex joint could be installed, finishing at 0600 h on 22 November. Once installed, the gooseneck and choke, booster, and kill moonpool hoses were made up, pressure tested, and installed by 1515 h. The landing and riser joints were also picked up, connected, and lowered to the moonpool, where work began to couple the riser tensioners to the landing ring, finishing by 0330 h on 23 November.

After the ROV was launched, a dummy-landing test 50 m from the well center was conducted successfully at 1730 h, and by 2030 h the lower marine riser package was landed on the lower BOP stack and locked in place. Riser running components were laid down by 0600 h on 24 November. The diverter assembly was made up and run into the hole to 2987 m DRF by 1700 h and then slowly pulled out of the hole to 2878 m DRF. The first cement plug was set at 2200 h and then pulled out of the hole to 2460 m DRF while waiting for cement. At 0445 h on 25 November, the BHA was run into the hole back to 2805 m DRF to confirm the top of cement and to begin the first cement plug pressure test. After confirming the cement plug, the BHA was pulled out of the hole to 2367 m DRF to set the second cement plug, pulled out of the hole to 2120 m DRF, and then pulled out of the hole to the surface by 1745 h.

The Hydralift Power Swivel (HPS) was parked and the riser running tool and riser guide head were installed, finishing at 2245 h. Once complete, the master bushing was removed while the ROV removed the hydrate build-up around the wellhead connector. The diverter and upper flex joint were picked up and laid down by 0400 h on 26 November, whereupon the BOP was disconnected, the wellhead was examined by the ROV, and the ROV was recovered to the surface by 0730 h.

The vessel moved to the LCA ~15 nmi from Hole C0002F and began recovering riser joints and the BOP at 1200 h on 27 November. The intermediate flex joint was recovered at 0200 h on 28 November, after which riser joint and BOP recovery continued. Recovery was briefly halted on 30 November to load the guide horn from the supply boat Shincho-Maru and resumed at 1000 h. The riser joints were all recovered and laid down by 0930 h on 1 December, after which the BOP was pulled out to the surface, landed on the BOP cart, and moved to its storage position aft of the moonpool by 1815 h.

Once the BOP was loaded on the cart, the HPS was rigged up again, and the vessel returned to Site C0002, arriving at 0630 h on 1 December. At Site C0002, the ROV dove at 0745 h, and the vessel shifted to the Hole C0002G long-term borehole monitoring system (LTBMS) so that the ROV could inspect the LTBMS. The vessel shifted back to Hole C0002F, and the ROV set the corrosion cap and checked the bull’s eyes, finishing by 1145 h. Once complete, the ROV began recovering and deploying transponders; 4 transponders were deployed and 10 were recovered, all completed by 0330 h on 2 December. The lower and middle guide horn sections were set on the BOP cart and connected by 0415 h; once connected, the vessel moved to Site C0012.

Hole C0002H

On 8 December 2012 at 0130 h, the ship moved 7 nmi north-northwest of Hole C0002H after preparations for coring began with the make up of the rotary core barrel (RCB) BHA. A short WOW period lasted until 1600 h, after which the vessel moved to within 5 nmi west of the well center. Drifting in began at 1800 h while the BHA was run into the hole, with a 1 h standby as an internal BOP ball valve malfunction was resolved. We continued to drift in to the well center, dropping the center bit at 0300 h on 9 December. The seafloor was tagged and confirmed at 1965 m DRF (1936.5 mbsl) at 0345 h, as indicated by an increase in weight on bit (WOB). We washed down the first 36 m then began drilling a 10⅝ inch hole at 0430 h, reaching 752 mbsf by 2000 h. The sinker bar was run down to recover the center bit for a wear check at 2145 h and dropped again at 2330 h, landing 13 min later. We then drilled to 1055 mbsf by 0730 h on 10 December. After a series of sweeping out the hole, drilling ahead, and sweeping again, the sinker bar was run at 1100 mbsf to recover the center bit in preparation for dropping the inner core barrel. At 1430 h, coring began, advancing 9.5 m to 1110 mbsf with recovery on deck at 1545 h. The BHA was stuck for the next 45 min, but after working and sweeping, the core barrel was dropped at 1845 h for the next coring advance. Coring to 1120 mbsf (9.5 m coring advance) began and the core barrel was recovered on deck at 2105 h. After the core barrel was recovered, the drill pipe was stuck in the hole again, but constant work freed the pipe at 0145 h on 11 December. The BHA was pulled out of the hole with reaming and laid down on deck by 0715 h. No obvious overpull or packoff indication was seen on the drill string, so preparations to return to RCB coring began. The vessel was moved 2 nmi from Hole C0002I for preparation for running the RCB BHA into Hole C0002I.

Hole C0002I

Hole C0002I operations began with drifting in at 1100 h on 11 December 2012 and dropping the center bit at 1630 h prior to spudding in at 1964.5 m DRF (1936 mbsl) 9 min later. The first 33.5 m was washed down and drilling ahead began from 1800 h, reaching 818.5 mbsf at 0600 h on 12 December. Another series of drilling ahead, sweeping, and hole cleaning began once past 905 mbsf. On 13 December, the bit reached 1360 mbsf at 0430 h, continuing until 0445 h on 13 December, when the 4S azimuth thruster shut down and dynamic positioning (DP) status changed to “yellow.” Once this occurred, the BHA was pulled out of the hole to 900 mbsf in preparation for emergency pulling out of the hole to the seafloor. The 4S azimuth thruster was restarted at 0600 h, and DP status returned to “green.” Circulation and hole cleaning began once green DP status was established and drilling back to 1105.5 mbsf began.

Drilling stopped at 1005 mbsf, whereupon the BHA became stuck at 1030 h on 13 December. Operations to free the stuck pipe began immediately. Attempts to recover the center bit at 1400 h on 14 December were unsuccessful; therefore, rig up of the Schlumberger wireline tool, the Free-Point Indicator Tool (FPIT), began at 1645 h. The stuck position was confirmed at 853 and 915 mbsf, after which the FPIT was rigged down for running the colliding tool to free the drill pipe at 0300 h on 15 December. The colliding tool was set and the explosive primed at 0700 h, when operations to install the colliding tool were suspended because of bad weather conditions, specifically high winds. The lower connection of the drill pipe on the rig floor was broken to run the colliding tool directly from the rig floor by 0745 h when the Schlumberger wireline winch failed. Troubleshooting began immediately and finished at 2130 h. The colliding tool was rigged up and the explosive was reset by 2315 h. The tool was run at 0300 h on 16 December, and the drill pipe was cut at 0345 h. The colliding tool was rigged down by 0515 h, and pulling out of the hole began. After the tools were laid down, the Chikyu moved to the LCA to load equipment and to perform maintenance on the HPS top drive, ending on 17 December. After pressure tests confirmed the integrity of the HPS repairs, we began making up and testing the RCB BHA for Hole C0002J.

Hole C0002J

The vessel moved to a position 3 nmi west of Hole C0002J, while the RCB BHA was made up and run into the hole from 1830 h on 17 December 2012. Spudding in Hole C0002J was confirmed at 1966 m DRF (1937.5 mbsl) at 0830 h on 18 December. The BHA was immediately jetted to 35 mbsf before drilling began. By 2330 h, the bit reached 872.5 mbsf before stopping to space out a single joint. Drilling continued until reaching 902 mbsf at 0045 h on 19 December. After sweeping out the hole, coring began at 0430 h. A total of seven RCB cores were collected, finishing at a total depth (TD) of 940 mbsf at 1930 h. Once coring was completed, kill mud was spotted and the RCB BHA was pulled out of the hole to the surface, and tool lay down was completed by 0545 h on 20 December. The vessel moved upstream 2 nmi to Hole C0002K and began preparations for coring in Hole C0002K.

Hole C0002K

Preparations for hydraulic piston coring system (HPCS)/extended shoe coring system (ESCS) coring started at 0545 h on 20 December 2012, and the BHA was run into the hole to 1966 m DRF (estimated water depth = 1937.5 mbsl) by 1645 h. The 11 inch BHA washed down the first 30 mbsf by 1700 h and then began drilling ahead to 200 mbsf. At 2230 h, coring began with the HPCS, ending after reaching 205.5 mbsf because of technical considerations. Switching to the extended punch coring system (EPCS) was completed by 0215 h on 21 December; four cores from 205.5 to 239 mbsf were cut by the EPCS. The low recovery of the last EPCS cores (n = 2) caused the switch to the ESCS from 239 mbsf. Five ESCS cores were collected from 239 to 286.5 mbsf, and the last core was cut at 2315 h on 21 December. Once coring was completed, the ESCS BHA was pulled out of the hole to 1780 m DRF by 0100 h on 22 December. Two stands of S-150 drill pipe were added, and then the BHA was run down to 1800 m DRF to WOW from 0300 to 0830 h on 22 December. Once the cold front had passed, the Chikyu was shifted to the Hole C0002L well center for more coring.

Hole C0002L

The ESCS BHA was run down from 1800 to 1960 m DRF by 1030 h, just above the seafloor. Spudding in Hole C0002L began at 1045 h (water depth = 1937.5 mbsl) and washing down proceeded to 42 mbsf, after which drilling to 277 mbsf was completed by 1830 h. ESCS coring began at 277 mbsf at 2000 h on 22 December 2012. Coring finished at a TD of 505.0 mbsf at 0215 h on 25 December. Once coring operations were completed, 35 m3 of kill mud was spotted in the hole and the BHA was pulled out of the hole to 1932 m DRF (above the seabed) by 0400 h. The BHA was laid down by 1030 h, after which the transponders were released and recovered (by 1230 h) by the watch boat, ending Hole C0002L operations. The Chikyu shifted to 2 nmi upstream of Site C0018 in preparation for LWD drilling.